# Natural Counterfactuals With Necessary Backtracking

### v2

@@ 4.2 Naturalness Constraints
To ensure an intervention’s feasibility, we stipulate that the resulting counterfactual value $an({\bf A})^{*}$ should fall within the data distribution, thus being natural relative to this distribution. We introduce a naturalness criterion, which confines the feasible support for $an({\bf A})^{*}$ in Eqn. 2 of the FIO framework.
Intuitively, the more frequently a value occurs, the more “natural” it is considered. Therefore, we assess this naturalness by examining the distribution characteristics, such as density, of each variable’s value ${\bf V}_{j}={\bf v}_{j}^{*}$ within ${\mathbf{AN}}({\bf A})$. This assessment is relative to the variable’s local mechanism $p({\bf V}_{j}|pa_{j}^{*})$, where  ${\bf V}_{j}\!\in\!\!{\mathbf{AN}}({\bf A})$ and $pa_{j}^{*}$ denotes the parent value of ${\bf V}_{j}$. Specifically, our approach determines whether the density of ${\bf v}_{j}^{*}$ surpasses a pre-established threshold $\epsilon$, utilized as a standard for appraising naturalness. [^2] The concept of “naturalness” can have various interpretations. In our context, it is defined by the data distribution in Structural Causal Models (SCMs). We do not claim that this is the only or uniquely best interpretation, but that it is a useful one to study and deploy.
@@ 4.2.1 Local Naturalness Criteria
Indeed, it might be difficult to find a universal definition for naturalness criteria, and thus, various criteria for naturalness can be explored. This allows for a more tailored approach to assessing naturalness in different contexts or applications. We start by assessing the naturalness of one variable’s value, ${\bf v}_{j}^{*}$, within the counterfactual data point $an({\bf A})^{*}$ in this section, followed by an examination of the overall naturalness of $an({\bf A})^{*}$ in the next section.
@@
We propose the following potential criteria based on entropy-normalized density, CDF of exogenous variables, and CDF of conditional distributions. The entropy-normalized naturalness measure evaluates the naturalness of ${\bf v}_{j}^{*}$ in relation to its local mechanism $p({\bf V}_{j}|{\mathbf{PA}}_{j}=pa_{j}^{*})$. The CDF-based measures, namely the latter two criteria, consider data points in the tails to be less natural. Each of these criteria has its own intuitive appeal, and their relative merits will be discussed subsequently. We formally establish the three criteria as follows:
@@
    Entropy-Normalized Measure:
    $p({\bf v}_{j}^{*}|pa_{j}^{*})e^{H({\bf V}_{j}|pa_{j}^{*})}>\epsilon$, where $H({\bf V}_{j}|pa_{j}^{*})=\mathbb{E}[-\log p({\bf V}_{j}|pa_{j}^{*})]$;
@@
    Exogenous CDF Measure: $\min(F({\bf u}_{j}^{*}),1-F({\bf u}_{j}^{*}))>\epsilon$, i.e., $\epsilon<F({\bf u}_{j}^{*})<1-\epsilon$;
@@
    Conditional CDF Measure: $\min(F({\bf v}_{j}^{*}|pa_{j}^{*}),1-F({\bf v}_{j}^{*}|pa_{j}^{*}))>\epsilon$, i.e., $\epsilon<F({\bf v}_{j}^{*}|pa_{j}^{*})<1-\epsilon$.
where the function $\min(\cdot)$ returns the minimum value between two given values.
@@
Choice (1): Entropy-Normalized Measure. Specifically, Choice (1), $p({\bf v}_{j}^{*}|pa_{j}^{*})e^{H({\bf V}_{j}|pa_{j}^{*})}$, can be rewritten as $e^{\log(p({\bf v}_{j}^{*}|pa_{j}^{*}))+\mathbb{E}[-\log p({\bf V}_{j}|pa_{j}^{*})]}$, where $-\log(p({\bf v}_{j}^{*}|pa_{j}^{*}))$ can be seen as the measure of surprise of ${\bf v}_{j}^{*}$ given $pa_{j}^{*}$ and $\mathbb{E}[-\log p({\bf V}_{j}|pa_{j}^{*})]$ can be considered as the expectation of surprise of the local mechanism $p({\bf V}_{j}|pa_{j}^{*})$ (Ash, 2012). Hence, the measure is the relative naturalness (i.e., negative surprise) of ${\bf V}_{j}$. Implementing this measure may be straightforward when employing a parametric SCM where the conditional distribution can be explicitly represented.
@@
Choice (2): Exogenous CDF Measure. If using a parametric SCM, we might directly measure differences on exogenous variables. However, in a non-parametric SCM, exogenous variables are not identifiable, and different noise variables may have different distributions. Hence, we choose to use the CDF of exogenous variables to align the naturalness of different distributions, based on a common assumption to achieve non-parametric SCMs in the machine learning system. Under the assumption, the support of the local mechanism $p({\bf V}_{j}|{\mathbf{PA}}_{j}=pa_{j}^{*})$ does not contain disjoint sets, and a nonlinear, non-parametric model ${\bf V}_{j}=f_{j}(pa_{j}^{*},{\bf U}_{j})$ is implemented, where $f$ is monotonically increasing with respect to ${\bf U}_{j}$. Noise ${\bf U}_{j}$ usually assumed to be a standard Gaussian (Lu et al., 2020). In this way, counterfactuals are identifiable, discussed in Sec. 4.4. Data points from the tails of a standard Gaussian can be thought of as less improbable events. Hence, ${\bf V}_{j}={\bf v}_{j}^{*}$ satisfies local $\epsilon$-natural generation when its exogenous CDF $F({\bf u}_{j}^{*})$ falls within the range $(\epsilon,1-\epsilon)$. In practice, for a single variable, ${\bf U}_{j}$ is a one-dimensional variable, and it is easier to enforce the measure than Choice (1), which involves conditional distributions.
@@
Choice (3): Conditional CDF Measure. The measure treats a particular value in the tails of local mechanism $p({\bf V}_{j}|pa_{j}^{*})$ as unnatural. Hence, ${\bf V}_{j}={\bf v}_{j}^{*}$ meets local $\epsilon$-natural generation when $F({\bf V}_{j}={\bf v}_{j}^{*}|pa_{j}^{*})$ falls within the range $(\epsilon,1-\epsilon)$ instead of tails. This measure can be used in parametric models where the conditional distribution can be explicitly represented. Notice, it can be easily used in non-parametric models when those models satisfy the assumption mentioned in Choice (2), and the measure is equivalent to Choice (2), since the CDF $F(X|pa_{j}^{*})$ has a one-to-one mapping with the CDF $F({\bf U}_{j})$, i.e., $F({\bf v}_{j}^{*}|pa_{j}^{*})=F({\bf u}_{j}^{*})$, when ${\bf v}_{j}^{*}=f(pa_{j}^{*},{\bf u}_{j}^{*})$.
4.2.2 $\epsilon$-Natural Generation
Based on the definition of local $\epsilon$-natural generation, we can define $\epsilon$-natural generation to judge whether the counterfactual value $an({\bf A})^{*}$ is natural.
@@ 4.2.2 $\epsilon$-Natural Generation
Given a SCM containing a set ${\bf A}$. A set ${\mathbf{AN}}({\bf A})$ contains all ancestors of ${\bf A}$ and ${\bf A}$ itself. Data point ${\mathbf{AN}}({\bf A})=an({\bf A})^{*}$ satisfies $\epsilon$-natural generation, if and only if, for any variable ${\bf V}_{j}\in{\mathbf{AN}}({\bf A})$, ${\bf V}_{j}={\bf v}_{j}^{*}$ satisfies local $\epsilon$-natural generation, where ${\bf v}_{j}^{*}$ is the value of ${\bf V}_{j}$ and $\epsilon$ is a small constant.
where the definition of local $\epsilon$-natural generation could be one of the optional definitions in 4.2.1 and a larger value of $\epsilon$ implies a higher standard for the naturalness of the generated data point $an({\bf A})^{*}$. To make intervention feasible, we require $an({\bf A})^{*}$ to meet $\epsilon$-natural generation.
4.3 Distance Measure for Necessary Backtracking
The naturalness constraints outlined in Eqn. 2 were comprehensively addressed previously. We now focus on defining the necessary backtracking distance in Eqn. 2 of the FIO framework to minimize the change in the counterfactual data point. Our analysis introduces two distinct definitions of distance in the counterfactual context. The first prioritizes minimizing changes in observable variables to minimize perceptible differences post-intervention. The second focuses on reducing alterations in local mechanisms, recognizing them as the inherent cost of intervention. Due to space limitations, we will primarily elaborate on minimal changes in observable variables. A more thorough investigation of changes in local mechanisms can be found in Sec. H.
Minimal Change in Observable Variables. The differences in perception are closely tied to the alterations in the values of observable (endogenous) variables resulting from the intervention. Intuitively, the total distance is the aggregate of distances across various variables in the ancestor set ${\mathbf{AN}}({\bf A})$. Therefore, we utilize the $L^{1}$ norm to measure this distance, as the $L^{1}$ norm offers the advantage of additive distances across variables. We define this distance as the perception distance:
$D(an({\bf A}),an({\bf A})^{*})=\left\|an({\bf A})-an({\bf A})^{*}\right\|_{1}$ (3)
@@ 4.3 Distance Measure for Necessary Backtracking
where $an({\bf A})$ and $an({\bf A})^{*}$ represent the actual value and counterfactual value of ${\bf A}$’s ancestors ${\mathbf{AN}}({\bf A})$ respectively, where ${\bf A}\in{\mathbf{AN}}({\bf A})$. Because endogenous variables may vary in scale, we utilize the standard deviation of variables to normalize each endogenous variable before computing the distance. This normalization ensures a consistent and fair evaluation of changes across all variables, irrespective of their individual scales.
@@
Implicitly, our distance metric favors changes in variables that are as proximal as possible to the target variable ${\bf A}$, since altering a more preceding variable typically results in changes to more downstream variables. When the value $an({\bf A})^{*}$, resulting from a hard intervention on ${\bf A}$, meets the $\epsilon$-natural generation criterion, it suggests that our distance metric $D(an({\bf A}),an({\bf A})^{*})$ becomes minimal, i.e., $|{\bf a}-{\bf a}^{*}|$. This effectively eliminates the need for backtracking in such cases. However, If a hard intervention does not meet the $\epsilon$-natural generation criterion, it becomes necessary to backtrack.
4.4 Identifiability of Natural Counterfactuals
In practice, we often do not know the form of structural equation models and noise variables are not identifiable from the observed variables. Hence, we assume causal structural models satisfy the conditions of the following theorem derived from Theorem $1$ of Lu et al. (2020) with proof in Sec. B, under which natural counterfactuals are identifiable with unknown structural equation models.
Theorem 4.1 (Identifiable Counterfactuals).
Suppose ${\bf V}_{i}$ satisfies the following structural causal model: ${\bf V}_{i}:=f_{i}({\mathbf{PA}}_{i},{\bf U}_{i})$ for any ${\bf V}_{i}\in{\bf V}$, where ${\bf U}_{i}\perp{\mathbf{PA}}_{i}$ and assume unknown $f_{i}$ is smooth and strictly monotonic w.r.t. ${\bf U}_{i}$ for fixed values of ${\mathbf{PA}}_{i}$. If we have evidence ${\bf E}={\bf e}$, with an intervention $do({\bf C}={\bf c}^{*})$, the counterfactual distribution of ${\bf B}$ is identifiable; that is, $p({\bf B}~{}|~{}do({\bf C}={\bf c}^{*}),{\bf E}={\bf e})$ is identifiable. [^3] If the assumption does not hold, identifiability may fail. For example, for $Y=XU_{1}+U_{2}$ where $Y$ and $X$ are endogenous variables and $U_{1}$ and $U_{2}$ are noise, the counterfactual outcome is not identifiable. Similarly, our approach can be applied to discrete variables satisfying the assumption of counterfactual stability developed by (Oberst & Sontag, 2019) in theory.
Following Feasible Intervention Optimization, we obtain the feasible intervention $do({\bf C}={\bf c}^{*})$. This theorem ensures the identifiability of our natural counterfactuals given $do({\bf C}={\bf c}^{*})$, under the stated assumptions. Note that the assumptions do not require more information on the structural functions or on the noise distributions. Hence, even without full information about the SCM, we can still use conditional distributions to infer natural counterfactuals, as well as non-backtracking counterfactuals, if the assumptions of the above theorem hold.
5 A Method for Generating Natural Counterfactuals
In this section, we provide a practical method for integrating natural counterfactuals into a machine learning system. In our scenario, we start with real-world data and a causal graph, lacking the SCM. We employ a non-parametric model to model the conditional distributions of endogenous variables and consider the model as a non-parametric SCM, making the assumption that each noise variable adheres to a standard Gaussian distribution. This learned SCM serves as the basis for generating natural counterfactuals.
To implement a particular method, we plug perception distance (Eqn. 3) and naturalness constraint (Choice (3) in Sec. 4.2.1) into Eqn. 2 of the FIO framework. Below is the equation of optimization:
$\underset{an({\bf A})^{*}}{\text{minimize}}\left\|an({\bf A})-an({\bf A})^{*}\right\|_{1}$ (4)

### v3

@@ 4.1 Naturalness Constraints
As indicated previously, the intended purpose of the naturalness constraint is to confine the counterfactual instance sufficiently within the support of the data distribution. Roughly and intuitively, the more frequently a value occurs, the more it is considered to be “natural”. Moreover, we would like to use a measure of naturalness that takes into account local mechanisms according to the given causal graph rather than just the joint distribution. Therefore, we propose to assess naturalness by examining the distribution characteristics, such as density, of each variable’s value ${\bf V}_{j}={\bf v}_{j}^{*}$ given the variable’s local mechanism $p({\bf V}_{j}|pa_{j}^{*})$, where  ${\bf V}_{j}\!\in\!\!{\mathbf{AN}}({\bf A})$ and $pa_{j}^{*}$ denotes its parental value setting. [^3] The notion of “naturalness” can have various interpretations. In our context, it is defined by the observed data distribution, as we assume we can only access observed data and intend the counterfactuals to be empirically supported. While this is obviously not the only plausible interpretation of naturalness, it is a useful one for our purpose.
@@ 4.1.1 Local Naturalness Criteria
We start by proposing some measures of the naturalness of one variable’s value, ${\bf v}_{j}^{*}$, within the counterfactual data point $an({\bf A})^{*}$ in this section, followed by defining a measure of the overall naturalness of $an({\bf A})^{*}$ in the next.
@@
We propose the following potential criteria based on entropy-normalized density, CDF of exogenous variables, and CDF of conditional distributions, respectively. The entropy-normalized naturalness measure evaluates the naturalness of ${\bf v}_{j}^{*}$ in relation to its local mechanism $p({\bf V}_{j}|{\mathbf{PA}}_{j}=pa_{j}^{*})$. The CDF-based measures, namely the latter two criteria, consider data points in the tails to be less natural. Each of these criteria has its own intuitive appeal, and their relative merits will be discussed subsequently. Below, we use $g_{l}({\bf v}_{j}^{*})$ to represent a (local) naturalness measure of ${\bf v}_{j}^{*}$. We consider the following three possible measures:
@@
    Entropy-Normalized Measure:$g_{l}({\bf v}_{j}^{*})=p({\bf v}_{j}^{*}|pa_{j}^{*})e^{H({\bf V}_{j}|pa_{j}^{*})}$, where $H({\bf V}_{j}|pa_{j}^{*})=\mathbb{E}[-\log p({\bf V}_{j}|pa_{j}^{*})]$;
@@
    Exogenous CDF Measure: $g_{l}({\bf v}_{j}^{*})=\min(F({\bf u}_{j}^{*}),1-F({\bf u}_{j}^{*}))$;
@@
    Conditional CDF Measure: $g_{l}({\bf v}_{j}^{*})=\min(F({\bf v}_{j}^{*}|pa_{j}^{*}),1-F({\bf v}_{j}^{*}|pa_{j}^{*}))$;
where the function $\min(\cdot)$ returns the minimum of the given values. When $g_{l}({\bf v}_{j}^{*})>\epsilon$, we say the ${\bf v}_{j}^{*}$ given its causal parents’ values satisfies the criterion of local $\epsilon$-natural generation.
Some comments on these choices are in order:
@@
Choice (1): Entropy-Normalized Measure. Specifically, Choice (1), $p({\bf v}_{j}^{*}|pa_{j}^{*})e^{H({\bf V}_{j}|pa_{j}^{*})}$, can be rewritten as $e^{\log p({\bf v}_{j}^{*}|pa_{j}^{*})+\mathbb{E}[-\log p({\bf V}_{j}|pa_{j}^{*})]}$, where $-\log p({\bf v}_{j}^{*}|pa_{j}^{*})$ can be seen as the measure of surprise of ${\bf v}_{j}^{*}$ given $pa_{j}^{*}$ and $\mathbb{E}[-\log p({\bf V}_{j}|pa_{j}^{*})]$ can be considered as the expectation of surprise of the local mechanism $p({\bf V}_{j}|pa_{j}^{*})$ (Ash, 2012). Hence, the measure quantifies the relative naturalness (i.e., negative surprise) of ${\bf V}_{j}$. Implementing this measure is usually straightforward when employing a parametric SCM where the conditional distributions can be explicitly represented.
@@
Choice (2): Exogenous CDF Measure. If using a parametric SCM, we might directly measure differences on exogenous variables. However, in a non-parametric SCM, exogenous variables are in general not identifiable, and different noise variables may have different distributions. Still, we may consider using the CDF of exogenous variables to align the naturalness of different distributions, based on a common assumption for non-parametric SCMs in the machine learning system. The assumption is that the support of the local mechanism $p({\bf V}_{j}|{\mathbf{PA}}_{j}=pa_{j}^{*})$ does not contain disjoint sets, the function $f_{j}$ in the SCM is monotonically increasing with respect to the noise variable ${\bf U}_{j}$, which is assumed to follow a standard Gaussian distribution (Lu et al., 2020). Data points from the tails of a standard Gaussian can be thought of as improbable events. Hence, ${\bf V}_{j}={\bf v}_{j}^{*}$ satisfies local $\epsilon$-natural generation when its exogenous CDF $F({\bf u}_{j}^{*})$ falls within the range $(\epsilon,1-\epsilon)$. In practice, for a single variable, ${\bf U}_{j}$ is a one-dimensional variable, and it is easier to enforce the measure than Choice (1), which involves conditional distributions.
@@
Choice (3): Conditional CDF Measure. The measure treats a particular value in the tails of local mechanism $p({\bf V}_{j}|pa_{j}^{*})$ as unnatural. Hence, ${\bf V}_{j}={\bf v}_{j}^{*}$ meets local $\epsilon$-natural generation when $F({\bf V}_{j}={\bf v}_{j}^{*}|pa_{j}^{*})$ falls within the range $(\epsilon,1-\epsilon)$ instead of tails. This measure can be used in parametric models where the conditional distribution can be explicitly represented. It can also be easily used in non-parametric models and the measure is equivalent to Choice (2) when those models satisfy the assumption mentioned in Choice (2), since the CDF $F({\bf V}_{j}|pa_{j}^{*})$ has a one-to-one mapping with the CDF $F({\bf U}_{j})$, i.e., $F({\bf v}_{j}^{*}|pa_{j}^{*})=F({\bf u}_{j}^{*})$, when ${\bf v}_{j}^{*}=f(pa_{j}^{*},{\bf u}_{j}^{*})$.
4.1.2 Global Naturalness Criteria
Given a local naturalness measure $g_{l}$, we simply define a global naturalness measure for $an({\bf A})^{*}$ as
$g_{n}(an({\bf A})^{*})=\min_{{\bf v}_{j}^{*}\in an({\bf A})^{*}}(g_{l}({\bf v}_{j}^{*})).$ (3)
That is, $g_{n}(an({\bf A})^{*})$ returns the smallest local naturalness value among members of ${\mathbf{AN}}({\bf A})$. Finally, we can define a criterion of $\epsilon$-natural generation to assess whether the counterfactual value $an({\bf A})^{*}$ is sufficiently natural.
@@ 4.1.2 Global Naturalness Criteria
Given a SCM containing a set ${\bf A}$, let ${\mathbf{AN}}({\bf A})$ contain all ancestors of ${\bf A}$ and ${\bf A}$ itself. A value setting ${\mathbf{AN}}({\bf A})=an({\bf A})^{*}$ satisfies $\epsilon$-natural generation, if and only if, $g_{n}(an({\bf A})^{*})>\epsilon$ and $\epsilon$ is a small constant.
Obviously, a larger value of $\epsilon$ implies a higher standard for the naturalness of the counterfactual value setting $an({\bf A})^{*}$. To consider only feasible interventions, we require $an({\bf A})^{*}$ to meet $\epsilon$-natural generation, which is a constraint used in FIO.
4.2 Distance Measure to Limit Backtracking
We now turn to the distance measure in Eqn. 2 of the FIO framework. We considered two distinct distance measures in this work. The first prioritizes minimizing changes in the observed causal ancestors of the target variable of the desired change. The second focuses on reducing alterations in local mechanisms, regarding them as inherent costs of an intervention. Due to space limitations, we will introduce here only the simpler measure in terms of minimal changes in the observable causal ancestors. A discussion of the other measure can be found in Sec. G.
For our purpose, the $L^{1}$ norm is a good choice, as it encourage sparse changes and thus sparse backtracking:
$D(an({\bf A}),an({\bf A})^{*})=\left\|an({\bf A})^{*}-an({\bf A})\right\|_{1}$ (4)
@@ 4.2 Distance Measure to Limit Backtracking
where $an({\bf A})$ and $an({\bf A})^{*}$ represent the actual value and counterfactual value of ${\bf A}$’s ancestors ${\mathbf{AN}}({\bf A})$ respectively, where ${\bf A}\in{\mathbf{AN}}({\bf A})$. Because endogenous variables may vary in scale, e.g., a normal distribution with a range of (,)($\infty$,$\infty$)( - ∞ , ∞ ) versus a uniform distribution over the interval $[0,1]$, we standardize each endogenous variable before computing the distance. This normalization ensures a consistent and fair evaluation of changes.
@@
Implicitly, this distance metric favors changes in variables that are proximal to ${\bf A}$, since altering a more remote variable typically results in changes to more downstream variables. In the extreme case, when the value $an({\bf A})^{*}$ corresponding to $do({\bf A}={\bf a}^{*})$, i.e., the one corresponding to the non-backtracking counterfactual, already meets the $\epsilon$-natural generation criterion, the distance metric $D(an({\bf A}),an({\bf A})^{*})$ will achieve the minimal value $|{\bf a}-{\bf a}^{*}|$. In such a case, no backtracking is needed and the non-backtracking counterfactual will be generated. However, If $do({\bf A}={\bf a}^{*})$ does not meet the $\epsilon$-natural generation criterion, it becomes necessary to backtrack.
4.3 Identifiablity of Natural Counterfactuals
As said, we assume we do not have prior knowledge of the functions of the SCM and so noise variables are in general not identifiable from the observed variables. However, if we assume the SCM satisfies the conditions of the following theorem, then the counterfactual instance resulting from a LBF intervention is identifiable.
Theorem 4.1 (Identifiable Natural Counterfactuals).
Given the causal graph and the joint distribution over ${\bf V}$, suppose ${\bf V}_{i}$ satisfies the following structural causal model: ${\bf V}_{i}:=f_{i}({\mathbf{PA}}_{i},{\bf U}_{i})$ for any ${\bf V}_{i}\in{\bf V}$, assume every $f_{i}$, though unknown, is smooth and strictly monotonic w.r.t. ${\bf U}_{i}$ for fixed values of ${\mathbf{PA}}_{i}$. Then, given an actual data point ${\bf V}={\bf v}$, with a LBF intervention $do({\bf C}={\bf c}^{*})$ (satisfying the criterion of $\epsilon$-natural generation), the counterfactual instance ${\bf V}={\bf v}^{*}$ is identifiable: ${\bf V}={\bf v}^{*}|do({\bf C}={\bf c}^{*}),{\bf V}={\bf v}$.
This theorem confirms the identifiability of our natural counterfactuals from the causal graph and the joint distribution over the observed variables. [^4] Given a substantial naturalness constraint, there may be one or more solutions for ${\bf C}$, or no solution at all. Which of these is the case can always be identified. The above theorem on identifiability of natural counterfactuals is concerned only with the cases in which a natural counterfactual exists, meaning that there is at least one solution for ${\bf C}$. Specifically, since $do({\bf C}={\bf c}^{*})$ satisfies the criterion of $\epsilon$-natural generation, it guarantees that the resulting counterfactual instance falls within the support of the observed joint distribution. Then, building on Theorem 1 from (Lu et al., 2020), we can demonstrate that using the actual data distribution allows for the inference of natural counterfactuals without knowing the functions or the noise distributions of the SCM.
5 A Practical Method for Generating Natural Counterfactuals
In this section, we provide a practical method for solving (approximately) the FIO problem described in the last section. We assume that we are given data sampled from the joint distribution of the endogenous variables and a causal graph, and that the underlying SCM satisfies the assumptions in Theorem 4.1. We learn a generative model for the endogenous variables from data, serving as an estimated SCM to generate natural counterfactuals: ${\bf V}_{i}:=\hat{f}_{i}({\mathbf{PA}}_{i},{\bf U}_{i})$ for $i=1,...,N$, where $\hat{f}_{i}$ is assumed to be reversible given ${\mathbf{PA}}_{i}$. Note that, unlike the functions in the true SCM, these learned functions in general do not generalize well to out-of-distribution data, as demonstrated in the experiments, which, recall, is a main motivation for employing natural counterfactuals instead. For simplicity, we assume the noise distribution is standard Gaussian, though the identifiability of natural counterfactuals only requires that noise variables are continuous and does not depend on the specific form of the noise distribution. The specific FIO problem we target plugs the distance measure (Eqn. 4) and the naturalness measure from Choice (3) in Sec. 4.1.1 (or Choice (2), which is equivalent to Choice (3) given the assumptions) into Eqn. 2:
$\underset{an({\bf A})^{*}}{\text{minimize}}\left\|an({\bf A})^{*}-an({\bf A})\right\|_{1}$ (5)